OpenStudy (anonymous):

1/3(24p-66)=3p+43 solve the equation

OpenStudy (anonymous):

(24p-66)/3=3p+43

OpenStudy (anonymous):

24p-66=(3p+43)3

OpenStudy (anonymous):

24p-66=9p+129

OpenStudy (anonymous):

24p-9p=129+66

OpenStudy (anonymous):

15p=195

OpenStudy (anonymous):

p=195/15

OpenStudy (anonymous):

p=13

OpenStudy (anonymous):

what happened to the

OpenStudy (anonymous):

\[\frac{ 1}{ 3 }\]

OpenStudy (anonymous):

in the beginning? @mustry

OpenStudy (anonymous):

1/3(24p-66)=1/3*24p-66*1/3=24p/3-66/3=(24p-66)/3

OpenStudy (anonymous):

u gt me?

OpenStudy (anonymous):

no sorry.. ;c

OpenStudy (anonymous):

let's do it step by step please? c:

OpenStudy (anonymous):

or draw it

OpenStudy (anonymous):